# ADO-LLM: Analog Design Bayesian Optimization with In-Context Learning of Large Language Models

### v1

@@ 1. Introduction
Large Language Models (LLMs) present great promise in addressing the above challenge and may be well-positioned to augment BO’s capabilities in analog circuit sizing tasks. We argue that the power of LLMs in circuit design originates from two key capabilities: the extensive prior knowledge embedded in their pre-trained data and their ability to perform in-context learning that incorporates human expertise. This allows LLMs to suggest design modifications and innovations with a high degree of relevance. Typically, the application of LLMs in this context follows an iterative loop, where previous design examples are demonstrated to the LLM, prompting it to generate new, potentially optimal design points. These points are then simulated, and the outcomes are used to further enrich the LLM’s training examples (Liu et al., 2024b). However, this few-shot generation process has limitations. The quality of optimization heavily depends on both the intrinsic capabilities of the LLM and the quality of the input demonstrations. Moreover, LLMs tend to produce solutions that closely mimic provided examples, showing a hesitance to extrapolate beyond the demonstrated data and explore new design areas.
@@ 3.1. Overview of the ADO-LLM Framework
Initially, the LLM agent is configured with circuit definition files and design specifications (see Figure 1 top-left). Then it utilizes a zero-shot approach to leverage its pre-trained knowledge for generating initial design points (see Figure 1 bottom-left). These points, enriched with embedded domain insights, initiate an iterative loop where design points are proposed, evaluated via the simulator, and used to update the dataset, thereby enhancing the learning context for subsequent iterations. The overall architecture and optimization flow is illustrated in Figure 1.
@@ 3.3.2. In-context Learning:
Circuit sizing problem is inherently a complex task that typically requires multiple steps of complex reasoning. In ADO-LLM, we follow the chain of thoughts (Wei et al., 2023) to prompt the LLM to decompose the problem into multiple steps:
@@
b. Balancing Trade-offs in the Design Specifications: We prompt the LLM to describe how to meet all design specifications for the interpreted circuit and explain the trade-offs between each objective. This step helps articulate the design knowledge pre-embedded within the training data.
@@
d. Injecting Human Expertise with Design Principles: We also utilize additional instruction termed “design principles” to improve the generation quality of the LLM agent. The adopted design principles instruct the LLM to ensure that all transistors operate within specific desired regions when proposing design points. These principles help the LLM utilize high-level simulator feedback effectively, while also preventing it from merely regurgitating trivial solutions from its training data.
@@ 4. Experiments
We compare ADO-LLM two baselines: the GP-BO model with random initialization and a single LLM proposer that uses zero-shot initialization to generate starting points. The number of initial design data is 5 for both methods. To ensure a fair comparison, the total number of evaluations of each method is set to be the same as that of ADO-LLM.
@@
Gain-Bandwidth Product (Spec) 1MHz
Gain (Spec) 60dB
Common-Mode Rejection Ratio (Spec) 75dB
Phase Margin (Spec) 60°
@@
Gain-Bandwidth Product (Norm) [0, 10MHz]
Gain (Norm) [0, 60dB]
Common-Mode Rejection Ratio (Norm) [0, 80dB]
Phase Margin (Norm) [0, 45°]
@@ 4.1. Two-Stage Differential Amplifier
We aim to design a circuit that satisfies the specifications of several performance metrics, including gain-bandwidth product (GBW), gain, common-mode rejection ratio (CMRR), phase margin (PM), and power consumption. These specifications are listed in LABEL:tab:ckt-amp2-info. To facilitate GP-BO for this multi-objective optimization task, we construct FOM as a linear combination of considered metrics:
(2) $FOM=\widetilde{GBW^{b}}+\widetilde{Gain^{b}}+\widetilde{CMRR^{b}}+\widetilde{PM^{b}}-\widetilde{Power^{b}}$
The superscript ${b}$ means the bounded value, i.e., a metric does not need to exceed beyond a corresponding specification bound. For metrics that contribute positively to FOM, i.e., GBW, Gain, CMRR, PM, the upper bound is defined as two times the maximum of the normalization range. The lower bound of the rest of the metrics is zero. The operator $\widetilde{\cdot}$ is a function to check whether the specification is satisfied, and then to normalize the metric. It is defined as
where the normalization range and the failed value of each metric are listed in LABEL:tab:ckt-amp2-info.
Results
For each method, we present its performance metrics, FOM, and number of missed specifications in Table 3. The best metric across all methods is bold, and a cross mark is appended to a metric when it fails to meet the corresponding design specification. Compared with GP-BO and LLM agent, ADO-LLM achieves the largest FOM, and even better than that acquired by GP-BO after 4x iterations of simulation. More importantly, only ADO-LLM satisfies all design specifications. While the baselines GO-BO and the LLM agent excel in optimizing several circuit performance metrics, they fail to balance these metrics in a limited simulation budget. This result demonstrates the data efficiency of ADO-LLMin analog circuit design tasks.
@@ Results
Gain (Spec) 60dB
Unit Gain Frequency (Spec) 10MHz
Hysteresis Error (Spec) 300mV
Voltage Offset (Spec) 20mV
@@
Gain (Norm) [0, 60dB]
Unit Gain Frequency (Norm) [0, 10MHz]
Hysteresis Error (Norm) [0, 300mV]
Voltage Offset (Norm) [0, 20mV]
@@
Voltage Offset (Failed) 40mV
@@
4.2. Hysteresis Comparator
Experimential Settings
(4) $FOM=\widetilde{Gain^{b}}+\widetilde{UGF^{b}}-\widetilde{V_{hys\_err}^{b}}-\widetilde{|V_{offset}^{b}|}-\widetilde{Power^{b}}$
LABEL:tab:ckt-comp-info presents our design specifications and configurations.
@@
Table 6. The best design of GP-BO with different initialization
@@ 4.3. The significance of Zero-Shot Initialization
Experimental Settings: We employed Gaussian Process-based Bayesian Optimization (GP-BO) for the design of the specified amplifier and comparator circuits. We compared two initialization strategies: uniform sampling from the entire search space and zero-shot initialization provided by the LLM proposer. We configured the experiments with 5 initial points, a batch size of 5 queries, and a total of 20 iterations.
@@ Results
Table 6 shows the best FOM acquired by GP-BO, where the abbreviation Amp2 stands for the two-stage amplifier, and Comp stands for hysteresis comparator. The initial data from LLM contributes to a better FOM in either circuit, and it is even comparable with that acquired after 80 iterations, as listed in Table 3 and Table 5. This result demonstrates that the domain-specific prior knowledge of LLM is able to provide good initialization parameter sets in potential high-value regions for analog circuit design tasks.
Table 7. The best design of LLM agent with different context examples
Method Circuit Sampler FOM # Missed Spec.
LLM agent w/o ICT Amp2 NA -1.78 3
LLM agent with ICT Amp2 Rand 5 -0.63 2
LLM agent with ICT Amp2 Top 5 0.20 1
LLM agent w/o ICT Comp NA -2.29 2
LLM agent with ICT Comp Rand 5 -2.53 3
LLM agent with ICT Comp Top 5 -1.35 1
@@ 4.4. The Effectiveness of In-Context Learning and Proposed Sampler
We design two variants of LLM agent: the first LLM agent is not provided with any demonstration. The second agent learns from few-shot examples that are uniformly sampled from the collected dataset. We compare these two variants with the proposed LLM agent which is demonstrated with few-shot examples with the highest top-k FOM. For all three variants, the number of sampled examples (if use examples) in each iteration is 5, and the total simulation budget is 105.
@@ Results
Table 7 reports the best design example of LLM agent variants in optimizing the amplifier and the comparator. Here the abbreviation ICT shorts for in-context learning. While random sampling works for the amplifier design task, this strategy degrades the zero-shot capacity of the LLM agent to optimize the comparator. With high-quality design parameters provided by the proposed sampler, the LLM agent achieves the best performance in both FOM and satisfaction of design specifications, demonstrating the effectiveness of the in-context learning and our sampling strategy for analog circuit design.
@@ 6. Limitation and Future Works
2. Optimality and Scalability of GP in BO: The Gaussian Processes (GP) used in our current BO setup effectively handle the design spaces demonstrated in this study; however, they may not scale efficiently to more complex or larger design challenges. In practical applications, as design spaces expand, the computational overhead and the inefficacy of optimizing all parameters simultaneously become significant constraints. A promising direction for future research is to implement hierarchical optimization strategies, where only a subset of parameters is optimized in each iteration, mimicking the approach of human experts. This could potentially increase both the efficiency and reliability of the design process. We leave these interesting directions to be explored in the future.

### v2

@@ 1. Introduction
Large Language Models (LLMs) present great promise in addressing the above challenge and may be well-positioned to augment BO’s capabilities in analog circuit sizing tasks. We argue that the power of LLMs in circuit design originates from two key capabilities: the extensive prior knowledge embedded in their pre-training data and their ability to perform in-context learning that can be utilized to incorporate domain expertise. This allows LLMs to suggest design modifications and innovations with a high degree of relevance. Typically, the application of LLMs in this context follows an iterative loop, where previous design examples are demonstrated to the LLM, prompting it to generate new, potentially optimal design points. These points are then simulated, and the outcomes are used to further enrich the LLM’s training examples (Liu et al., 2024b). However, this few-shot generation process has limitations. The quality of optimization heavily depends on both the intrinsic capabilities of the LLM and the quality of the input demonstrations. Moreover, LLMs tend to produce solutions that closely mimic the provided examples, showing a hesitance to extrapolate beyond the demonstrated examples and explore new design areas.
@@ 3.1. Overview of the ADO-LLM Framework
Initially, the LLM agent is configured with circuit definition files and design specifications (see Figure 1 top-left). Then it utilizes a zero-shot approach to leverage its relevant analog design knowledge obtained from the vast amount of model pre-training data for generating initial design points (see Figure 1 bottom-left). These points, enriched with embedded domain insights, initiate an iterative loop where design points are proposed, evaluated via the simulator, and used to update the dataset, thereby enhancing the learning context for subsequent iterations. The overall architecture and optimization flow is illustrated in Figure 1.
@@ 3.3.2. In-context Learning:
Circuit sizing is inherently a complex task that typically requires multiple steps of reasoning. In ADO-LLM, we follow the chain of thoughts (Wei et al., 2023) to prompt the LLM to decompose the problem into multiple steps:
@@
b. Balancing Trade-offs in the Design Specifications: We prompt the LLM to describe how to meet all design specifications for the interpreted circuit and explain the trade-offs between each objective. This step helps articulate the design knowledge pre-embedded within the large language dataset used for training of LLMs.
@@
d. Injecting Human Expertise with Design Principles: We also utilize additional instructions termed “design principles” to improve the generation quality of the LLM agent. The adopted design principles instruct the LLM to ensure that all transistors operate within specific desired regions when proposing design points. These principles help the LLM utilize high-level simulator feedback effectively, while also preventing it from merely regurgitating trivial solutions from its training data.
@@ 4. Experiments
We compare ADO-LLM with two baselines: the GP-BO model with random initialization and a single LLM proposer that uses zero-shot initialization to generate starting points. The number of initial design data is 5 for both methods. To ensure a fair comparison, the total number of evaluations of each method is set to be the same as that of ADO-LLM.
@@
Gain-Bandwidth Product (Spec) $\geq$ 1MHz
Gain (Spec) $\geq$ 60dB
Common-Mode Rejection Ratio (Spec) $\geq$ 75dB
Phase Margin (Spec) $\geq$ 60°
@@
Gain-Bandwidth Product (Norm) [0MHz, 10MHz]
Gain (Norm) [0dB, 60dB]
Common-Mode Rejection Ratio (Norm) [0dB, 80dB]
Phase Margin (Norm) [0°, 45°]
@@ 4.1. Two-Stage Differential Amplifier
We aim to design a circuit that satisfies the specifications of several performance metrics, including gain-bandwidth product (GBW), gain, common-mode rejection ratio (CMRR), phase margin (PM), and power consumption. These specifications are listed in LABEL:tab:ckt-amp2-info. To facilitate GP-BO for this multi-objective optimization task, we construct FOM as a linear combination of the considered metrics:
@@
(2) $FOM=\widetilde{GBW}^{b}+\widetilde{Gain}^{b}+\widetilde{CMRR}^{b}+\widetilde{PM}^{b}-\widetilde{Power}$
The operator $\widetilde{\cdot}$ is a function to check whether the specification is satisfied, and then to normalize the metric. It is defined as
for a metric $m$ where the normalization range and the failed value of each metric are listed in LABEL:tab:ckt-amp2-info.
The superscript ${b}$ means the bounded value, i.e., a metric does not need to exceed far beyond its design specification. This bound is set to 2 for the normalized values of GBW, Gain, CMRR, and PM.
Results
For each method, we present its performance metrics, FOM, and number of missed specifications in Table 3. The best metric across all methods is bold, and a cross mark is appended to a metric when it fails to meet the corresponding design specification. Compared with GP-BO and LLM agent, ADO-LLM achieves the largest FOM, which is even better than that acquired by GP-BO after 4x iterations of simulation. More importantly, only ADO-LLM satisfies all design specifications. While the baselines GO-BO and the LLM agent excel in optimizing several circuit performance metrics, they fail to balance these metrics with a limited simulation budget. This result demonstrates the data efficiency of ADO-LLM in analog circuit design tasks.
4.2. Hysteresis Comparator
Experimential Settings
@@ 4.2. Hysteresis Comparator
Gain (Spec) $\geq$ 60dB
Unit Gain Frequency (Spec) $\geq$ 10MHz
Absolute Hysteresis Error (Spec) $\leq$ 300mV
Absolute Voltage Offset (Spec) $\leq$ 20mV
@@
Gain (Norm) [0dB, 60dB]
Unit Gain Frequency (Norm) [0mHz, 10MHz]
Hysteresis Error (Norm) [0mV, 300mV]
Absolute Voltage Offset (Norm) [0mV, 20mV]
@@
Absolute Voltage Offset (Failed) 40mV
@@
(4) $FOM=\widetilde{Gain}^{b}+\widetilde{UGF}^{b}-\widetilde{V_{hys\_err}}-\widetilde{|V_{offset}|}-\widetilde{Power}$
LABEL:tab:ckt-comp-info presents our design specifications and configurations. The upper bound is set to 2 for the normalized values of Gain and UGF in our experiments.
@@ Results
Table 6. The best design of GP-BO with different initializations
@@ 4.3. The significance of Zero-Shot Initialization
We employed Gaussian Process-based Bayesian Optimization (GP-BO) for the design of the specified amplifier and comparator circuits. We compared two initialization strategies: uniform sampling from the entire search space and zero-shot initialization provided by the LLM proposer. We configured the experiments with 5 initial points, a batch size of 5 queries, and a total of 20 iterations.
@@ Results
Table 6 shows the best FOM acquired by GP-BO, where the abbreviation Amp2 stands for the two-stage amplifier, and Comp stands for hysteresis comparator. The initial data from the LLM contributes to a better FOM in either circuit, which is even comparable with that acquired after 80 iterations, as listed in Table 3 and Table 5. This result demonstrates that the domain-specific prior knowledge of the LLM is able to provide good initialization parameter sets in potential high-value regions for analog circuit design tasks.
@@ 4.4. The Effectiveness of In-Context Learning and Proposed Sampler
Table 7. The best design of the LLM agent with different in-context examples
Method Circuit Sampler FOM # Missed Spec.
LLM agent w/o ICL Amp2 NA -1.78 3
LLM agent with ICL Amp2 Rand 5 -0.63 2
LLM agent with ICL Amp2 Top 5 0.20 1
LLM agent w/o ICL Comp NA -2.29 2
LLM agent with ICL Comp Rand 5 -2.53 3
LLM agent with ICL Comp Top 5 -1.35 1
@@
We design two variants of the LLM agent: the first LLM agent is not provided with any demonstration. The second agent learns from few-shot demonstrations that are uniformly sampled from the collected dataset. We compare these two variants with the proposed LLM agent which is provided with few-shot demonstrations with the highest top-k FOM. For all three variants, the number of sampled demonstrations if there is any in each iteration is 5, and the total simulation evaluation budget is 105.
@@ Results
Table 7 reports the best design example of the LLM agent variants in optimizing the amplifier and the comparator. Here the abbreviation ICL stands for in-context learning. While using random sampling to provide demonstrations improves the performance of the amplifier circuit, this strategy degrades the zero-shot capacity of the LLM agent when optimizing the comparator. With the high-quality design demonstrations provided by the proposed sampler, the LLM agent achieves superior performance in both the FOM and the number of specifications met, demonstrating the effectiveness of the in-context learning and our sampling strategy for analog circuit design.
@@ 6. Limitation and Future Works
2. Optimality and Scalability of GP in BO: The Gaussian Processes (GP) used in our current Bayesian Optimization (BO) setup effectively manage the design spaces of the circuits examined in this study. However, they may not scale well for more complex designs. In practical applications, as design spaces expand, the computational overhead and the inefficacy of optimizing all parameters simultaneously become significant constraints. A promising direction for future research is to implement hierarchical optimization strategies, where only a subset of parameters is optimized in each iteration, mimicking the approach of human experts. This could potentially increase both the efficiency and reliability of the design process. We leave these interesting directions to be explored in the future.